(d) Suppose Mary purchased a pension scheme with a principal of $3,500,000 and the annual interest rate is 8%. How much will she receive per quarter, forever? (3 marks)

(e) Suppose Mary has $50,000 in the bank account earning interest of 2% payable annually. As she plans to go travel for a month in Europe, she can save an additional $3,500 per month. How much will she accumulate 6 years from now? (correct to 2 decimal places)
Answer #1

(d) Suppose Mary purchased a pension scheme with a principal of $3,500,000 and the annual interest rate is 8%. How much will she receive per quarter, forever?

Interest rate per quarter = 8%/4 = 2%

The amount received per quarter = Principal x Interest rate per quarter

The amount received per quarter = 3,500,000 * 0.02

The amount received per quarter = $70,000

(e) Suppose Mary has $50,000 in the bank account earning interest of 2% payable annually. As she plans to go travel for a month in Europe, she can save an additional $3,500 per month. How much will she accumulate 6 years from now?

Monthly rate = 2%/12 = 0.001666666667

The future value of the annuity:

n = 6 * 12 = 72

PMT FV = =*[(1 + r) - 1

3,500 FV =- 0,001666666667 * ((1+0.001666666667)72_11

FV = 2,099,999.99958 * 0.1273842327

FV = $267,506.888616499

The future value of $50,000

FV = PV * (1 + r)^n

FV = 50,000 * (1 + 0.02)^6

FV = $56,308.1209632

Total accumulated amount in 6 years = 267,506.888616499 + 56,308.1209632

Total accumulated amount in 6 years = $323,815.00958

Can you please upvote? Thank You :-)